OpenStudy (anonymous):

Is there anyone in connexus? I need help with Algebra 1

OpenStudy (anonymous):

im in 8grade

OpenStudy (anonymous):

me too whats the question

OpenStudy (black_ninja123):

I can help you but im not in connections

OpenStudy (stuck-help):

im in connexux honors algebra 1

OpenStudy (anonymous):

11. x+y=8 and y=3x go u know how to solve this?

OpenStudy (anonymous):

do*

OpenStudy (stuck-help):

i dont sorry

OpenStudy (anonymous):

yea i can do it, can u help with mine? have u taken the semester 2 unit 4 test?

OpenStudy (anonymous):

The Linear test?

OpenStudy (anonymous):

unit 4 lesson 10

OpenStudy (anonymous):

nope imma still on the portfolio :/

OpenStudy (anonymous):

what lesson

OpenStudy (anonymous):

Systems of Equations and Inequalities Portfolio

OpenStudy (anonymous):

well not the portfolio part that imma doing im doing the sample work

OpenStudy (stuck-help):

i think i might have that answer then

OpenStudy (anonymous):

x+y=8 and y=3x 12. … 13. X+3=y and 3x+4y=7 14. …. 15. Y=-2x+6 and 3y-x+3=0 16. …. 17. y-2x=3 and 3x-2y=5 18. …. 19. 2=2y-x and 23=5y-4x 20. … 21. 7x-2y=1 and 2y=x-1 this is what imma stuck on

OpenStudy (stuck-help):

do you know how to get the answer sheets i dont have it that was semester A for me

OpenStudy (anonymous):

no i dont

OpenStudy (stuck-help):

go into the recourse backpack,recourse packets,what unit your in download the link

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Damn thx i didnt know that :D

OpenStudy (stuck-help):

yeah but most of the time you have to have to show your work and that only has the answers

OpenStudy (anonymous):

are u in FLVS? algebra 1? what lesson r u on?

OpenStudy (anonymous):

heyitme i have it thx tho